Sally travels 980m south 430m west find the displacement and distance?
1
Expert's answer
2020-03-11T14:11:55-0400

Distance:


"d=430+980=1410\ m"

Displacement:


"s=\sqrt{430^2+980^2}=1070\ m"

Direction:


"\theta=\arctan{\frac{430}{980}}=23.7\degree\text{W of S}"